OpenStudy (anonymous):

Show that the equation is not an identity by finding a value of x for which both sides are defined but not equal. cos x - cos x sinx = cos^3 x A. π B. 0 C. π/4 D. π/2

OpenStudy (campbell_st):

start by dividing each term by cos(x) 1 - sin(x) = cos^2(x) and cos^2(x) = 1 - sin^2 (x) from the trig identity sin^2 + cos^2 = 1 so you have 1 - sin(x) = 1 - sin^2(x) or sin^2(x) - sin(x) = 0 solve for x by factorising the quadratic equation.